Analyze a Diesel cycle using a constant specific heat air standard model (cP = 1.004 kJ/kg-K, R = 0.287 kJ/kg-K) with the following properties: The temperature and pressure at the beginning of the compression stroke are 300 K and 100 kPa, the compression ratio is 22, and the specific heat addition is 800 kJ/kg. Find the thermal efficiency, cut off ratio, maximum temperature in the cycle, and maximum pressure in the cycle.
Answer #1

T= 300k , P = 100 kPa compression ratio (Y) = 22 Heat Addition = 800 ICT 14 I. (Vok cp-c = R a = 1.004 - 0.287 a = 0.717 K =Ta = Tix 8X 1s = cut off ratio] 1.4 To = 300 x 1.7414t = 667.982 n Heat Rejection = ax [Te-T,} = 0.714 [667.982-300) • = 26